melissa's has a pickle that contains 65 coins and worth $8.65.The pickle jar only has nickles and quarters let q represent q and n represent nickles.(A:what is the system that models this situation.(B:which of these is a solution to the following system 29 quarters and 36 nickles 27 quarters and 38 nickles 26 quarters and 39 nickles

Answer by addingup(3677)

n+q = 65 subtract q from both sides:
n = 65-q
:
0.05n+0.25q = 8.65 substitute for n:
0.05(65-q)+0.25q = 8.65
3.25-0.05q+0.25q = 8.65
0.20q = 5.40
q = 27 the pickle jar has 27 quarters and (65-27=38) 38 nickels
----------
Check
27*0.25+38*0.05 = 8.65 Correct.
You do the rest.